OpenStudy (anonymous):

So I answered 5 more with the help but I don't understand this cause I believe it's a decimal... what 2 numbers multiply to -7 and add to be 6?

OpenStudy (campbell_st):

ok...can you list the factors of 7...?

OpenStudy (campbell_st):

7 is prime so there are only 2 factors

OpenStudy (anonymous):

1 and 7

OpenStudy (anonymous):

So that's my answer. -1 and 7...

OpenStudy (anonymous):

Thank you campbell_st

OpenStudy (campbell_st):

great if you do some multiplication and addition 1 x 7 = 7 1 + 7 = 8 -1 x -7 = 7 -1 + -7 = -8 -1 x 7 = -7 -1 + 7 = 6 <====== the correct pairing 1 x -7 = -7 1 + -7 = -6 so looking at what is above... if the factors have the same sign, the product is positive 1 factor negative then the product is negative. The key to the correct solution is looking at the sum... if the sum is positive then the larger factor is positive if the sum is negative, the larger factor is negative. hope it makes sense

OpenStudy (anonymous):

That's so helpful!!